40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Message ID: 97     Entry time: Sun Aug 21 17:16:03 2016
Author: Gabriele 
Type: General 
Category: Daily Progress 
Subject: Ring down measurement with new disk 

After fixing the setup, the measurement with the new disk looks great. After more than one hour the first two modes are still ringing down, meaning that the Q's are larger than 10 millions.

Here's the comparison of the spectrum before and after the excitation, with the identified modes:

I used about 4200 s of data to fit the ring downs. Most fits are good. In a couple of cases the peak splitting is large and the algorithm fails to fit the beats:

In summary, here are the Q values for all modes. Despite not being annealed, this disk shows very large Q's

ELOG V3.1.3-